at a certain concert, 31% of the audience was under 20 years old. A random sample of n= 298 members of the audience was selected. the the Ui, the mean of the distribution of sample proportions. Write only a number as your answer. Do not write as a percentage.
Answer #1

Solution

Given that,

p = 0.31

1 - p = 1 -0.31 =0.69

n = 298

mean = = p =0.31